# Equations of genus $4$ curves from their theta constants

### v2

@@ 1. Introduction
In genus 4 we are faced with a different situation, as the moduli space of genus $4$ curves with a full level $2$ structure is not unirational. (This will be proved in a later article [HPS24a].) This fact may explain why the general problem has remained open, only garnering results in special cases. For instance, Schottky [Sch88] managed to solve the problem in the case where one of the theta constants vanishes, which is equivalent to the canonical quadric being a cone.
@@
Recall from [FR70, Chapter 5] that the Schottky modular form is related to the Prym construction as follows. Given an étale double cover $\tilde{C}\to C$, the Prym variety $\operatorname{Prym}(\tilde{C}/C)=\ker(\text{Nm}:\operatorname{Jac}(\tilde{C})\rightarrow\operatorname{Jac}(C))^{\circ}$ is in a natural way a principally polarized abelian variety. In particular, the theta constants of $\operatorname{Prym}(\tilde{C}/C)$ must satisfy the quartic Riemann identities. The Schottky modular form is derived from these identities by expressing the theta constants of $\operatorname{Prym}(\tilde{C}/C)$ in terms of the theta constants of $\operatorname{Jac}(C)$.
@@ 1.1. Motivation and applications
In addition to the intrinsic value of recovering a genus 4 curve from its theta constants (or its period matrix), our method also has several applications. It is interesting in arithmetic and complex geometry to be able to find higher genus curves with special properties. Being able to reconstruct a curve from the theta constants (or period matrix) of its Jacobian allows us to construct Jacobians with interesting properties and find corresponding curves that inherit those properties. Below we list a couple of examples that illustrate this strategy. These examples were computed using our Magma [BCP97] implementation of our algorithm [HPS24c].
@@
  • Constructing modular abelian varieties. Shimura associates to a Hecke newform $f$ with level $\Gamma_{1}(N)$ a subvariety $A_{f}$ of the modular Jacobian $J_{1}(N)$. This $A_{f}$ is simple and $\operatorname{End}_{\mathbb{Q}}(A_{f})\otimes\mathbb{Q}$ contains the field $K$ generated by the coefficients of the $q$-expansion of $f$. Furthermore, $K$ is totally real and one has $[K:\mathbb{Q}]=\dim(A_{f})$. Conversely, Khare and Wintenberger [KW09a, Corollary 10.2], [KW09b] have shown that every simple abelian variety $A$ over $\mathbb{Q}$ with the property that $\operatorname{End}_{\mathbb{Q}}(A)\otimes\mathbb{Q}$ contains an RM field $K$ such that $[K:\mathbb{Q}]=\dim(A)$, must be modular. This means that there exists a Hecke newform $f$ with level $\Gamma_{1}(N)$ such that $A$ is isogenous to $A_{f}$. Furthermore, $N$ is then equal to the conductor of $A$.
@@ 1.2. Previous work
[…]
@@
rather than the map $\varphi$. The novelty of the present paper is a link between the map $\varphi$ and the Prym construction. This allows us prove that the knowledge of $\varphi$ for just one two-torsion point $\eta$ is sufficient to recover the curve $C$. In this way our method is more efficient than Lehavi’s, but our result does not logically imply his main theorem. The reason is that we do not only use the tritangent planes, but in addition information coming from the Prym variety.
@@ 1.3. Structure of the article
The article is organized as follows. In Section 2 we begin by defining the objects that will play key roles in our construction. In Section 3 we describe our main result: an algorithm that takes theta constants of the Jacobian of a genus 4 curve as its input and, using only elementary arithmetic, extraction of square roots, and solving linear systems of equations, outputs the corresponding quadric and cubic whose solution set is equal to the image of the canonical embedding of the corresponding curve. In Section 4 we present two examples as an application of our method. First, we construct a gluing with an interesting endomorphism algebra, and then the curve whose Jacobian is the modular abelian variety discussed above, i.e., with real multiplication by the maximal real subfield of $\mathbb{Q}(\zeta_{15})$. Finally, in Section 5 we briefly discuss how our method can be applied to find more interesting curves, as well as how it can be generalized to higher genus.
@@ 1.4. Acknowledgements
The authors would like to thank Thomas Bouchet, Nils Bruin, Edgar Costa, Igor Dolgachev, Noam Elkies, Gavril Farkas, Sam Grushevsky, Avi Kulkarni, Elisa Lorenzo García, David Lubicz, Martin Raum, and Christophe Ritzenthaler for helpful conversations during the writing of this article. Furthermore, we thank Brendan Hassett, David Lehavi, and Jeroen Sijsling for reading a previous version of the manuscript and giving suggestions to improve the article.
@@ 2.1. The Prym canonical map
[…]
@@ 2.2. Two linear maps
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 3.2. Finding the quadric
We will now construct an example over $\mathbb{F}_{37}$ and verify the required properties with the aid of Magma. (See the files Lemma-Section-3.2-proofs.m and construct-Lemma-Section-3.2-proofs.m in [HPS24c].) Consider the plane quartic $X\subset\mathbb{P}^{2}$ with equation
@@ 3.3. Finding a cubic
[…]
@@
[…]
@@ 3.4. Main result
1 Choose the two-torsion point given by $\eta=\frac{1}{2}\begin{bmatrix}0&0&0&0\\ 1&0&0&0\end{bmatrix}$.
2 Compute the theta constants for the plane quartic $X$ via the Schottky-Jung relations
@@
Although our main result Theorem 3.11 requires the base field $k$ to be algebraically closed, with additional information one can recover equations over arithmetic base fields. Suppose we are given the theta constants of a curve $C$ as complex numbers, but we expect that $C$ is defined over a number field $K$. If in addition to these theta constants, we are also given a big ($g\times 2g$) period matrix for $C$ with respect to a $K$-rational basis for $\operatorname{H}^{0}(C,\Omega_{C})$, we can recover equations for $C$ defined over $K$ as follows. First applying our theorem, we obtain approximate equations for $C$ with coefficients in $\operatorname{\mathbb{C}}$. Denote by $\widetilde{C}$ this complex curve, which is isomorphic to $C$ over $\operatorname{\mathbb{C}}$.
@@
When using the Schottky-Jung relations one takes a square root to compute the $\vartheta_{X}\big{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\big{]}$. This leads to a sign choice that has to be made correctly. For this one can use the discussion following [Gla80, Theorem 3.1] where Glass makes use of the quartic Riemann identities to find a correct choice of signs. At this point, the assumption that the $\vartheta_{C}\big{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\big{]}$ are the theta constants of a genus $4$ curve is used. Indeed, the latter implies that the Schottky modular form vanishes. By construction of this modular form, this is equivalent to requiring that the $\vartheta_{X}\big{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\big{]}$ satisfy the quartic Riemann identities.
@@ 4. Examples
To remedy this, we call the command SomePrincipalPolarizations from the ModularCurves GitHub repository [${}^{+}$23]. This produces several big period matrix candidates. Evaluating the Schottky modular form on the first of these yields a complex number with absolute value $6.2124\cdot 10^{-300}$ when calculated with precision 300, indicating that this principally polarized abelian variety is likely a Jacobian. Applying our method then produces the corresponding genus 4 curve, which, after a change of variable, is isomorphic to the curve $C$ in $\mathbb{P}^{3}$ given by the following equations.
@@ 5. Future work
To rule out CM fields containing primitive roots of unity, Moonen and Oort [MO11] defined the notion of a Weyl CM field. These are CM fields $K$ of degree $[K:\mathbb{Q}]=2g$ such that the Galois group of the splitting field is as large as possible, i.e., $(\mathbb{Z}/2)^{g}\rtimes S_{g}$. They put forward the conjecture that for a given fixed $g\geqslant 4$ there are finitely curves of genus $g$ whose Jacobian has CM by a Weyl CM field.
@@ 5.2. Higher genus
In work in progress [HPS24b] we plan to generalize the Milne correspondence to arbitrary $g$. This would give a bijection between pairs of odd theta hyperplanes in $\mathbb{P}^{g-1}$ for $C$ and odd theta hyperplanes in $\mathbb{P}^{g-2}$ for the Prym variety $\operatorname{Prym}(C,\eta)$. Notice that in this context the Prym variety is no longer automatically a Jacobian. Nevertheless, one can define odd theta hyperplanes for arbitrary p.p.a.v.s; however, these will lack the geometric interpretation as multitangents of a canonically embedded curve (unless the p.p.a.v. happens to be a Jacobian).
@@
[…]

### v3

@@ 1. Introduction
In genus 4 we are faced with a different situation, as the moduli space of genus $4$ curves with a full level $2$ structure is not unirational. (This will be shown in future work.) This fact may explain why the general problem has remained open, only garnering results in special cases. For instance, Schottky [Sch88] managed to solve the problem in the case where one of the theta constants vanishes, which is equivalent to the canonical quadric being a cone.
@@
Recall from [FR70, Chapter 5] that the Schottky modular form is related to the Prym construction as follows. Given an étale double cover $\tilde{C}\to C$, the Prym variety $\operatorname{Prym}(\tilde{C}/C)=\ker(\text{Nm}:\operatorname{Jac}(\tilde{C})\rightarrow\operatorname{Jac}(C))^{\circ}$ is in a natural way a principally polarized abelian variety. In particular, the theta constants of $\operatorname{Prym}(\tilde{C}/C)$ must satisfy the quartic Riemann theta formulas. The Schottky modular form is derived from these identities by expressing the theta constants of $\operatorname{Prym}(\tilde{C}/C)$ in terms of the theta constants of $\operatorname{Jac}(C)$.
@@ 1.1. Motivation and applications
In addition to the intrinsic value of recovering a genus 4 curve from its theta constants (or its period matrix), our method also has several applications. It is interesting in arithmetic and complex geometry to be able to find higher genus curves with special properties. Being able to reconstruct a curve from the theta constants (or period matrix) of its Jacobian allows us to construct Jacobians with interesting properties and find corresponding curves that inherit those properties. Below we list a couple of examples that illustrate this strategy. These examples were computed using our Magma [BCP97] implementation of our algorithm [HPS24b].
@@
  • Constructing modular abelian varieties. Shimura associates to a Hecke newform $f$ with level $\Gamma_{1}(N)$ a subvariety $A_{f}$ of the modular Jacobian $J_{1}(N)$. This $A_{f}$ is simple and $\operatorname{End}_{\mathbb{Q}}(A_{f})\otimes\mathbb{Q}$ contains the field $K$ generated by the coefficients of the $q$-expansion of $f$. Furthermore, $K$ is either totally real or a CM field and one has $[K:\mathbb{Q}]=\dim(A_{f})$. Conversely, Khare and Wintenberger [KW09a, Corollary 10.2], [KW09b] have shown that every simple abelian variety $A$ over $\mathbb{Q}$ with the property that $\operatorname{End}_{\mathbb{Q}}(A)\otimes\mathbb{Q}$ contains a field $K$ such that $[K:\mathbb{Q}]=\dim(A)$, must be modular. This means that there exists a Hecke eigenform $f$ with level $\Gamma_{1}(N)$ such that $A$ is isogenous to $A_{f}$. Furthermore, $N$ can be taken to be $M^{\dim(A)}$ where $M$ is the conductor of $A$.
@@ 1.2. Previous work
${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))\times_{\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}S^{2}\operatorname{H}^{0}(\Omega_{C})}$${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))}$${S^{2}\operatorname{H}^{0}(\Omega_{C})}$${\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}$$\scriptstyle{\varphi}$ (1.1)
@@
rather than the map $\varphi$. The novelty of the present paper is a link between the map $\varphi$ and the Prym construction. This allows us prove that the knowledge of $\varphi$ for just one two-torsion point $\eta$ is sufficient to recover the curve $C$. In this way our method is more efficient than Lehavi’s, but our result does not logically imply his main theorem. The reason is that we use not only the tritangent planes, but also information coming from the Prym variety.
@@ 1.3. Structure of the article
The article is organized as follows. In Section 2 we begin by defining the objects that will play key roles in our construction. In Section 3 we describe our main result: formulas for the equations of a genus 4 curve in terms of the theta constants of its Jacobian. In Section 4 we present two examples as an application of our method. First, we construct a gluing with an interesting endomorphism algebra, and then the curve whose Jacobian is the modular abelian variety discussed above, i.e., with real multiplication by the maximal real subfield of $\mathbb{Q}(\zeta_{15})$. Finally, in Section 5 we briefly discuss how our method can be applied to find more interesting curves, as well as how it can be generalized to higher genus.
@@ 1.4. Acknowledgements
The authors would like to thank Thomas Bouchet, Nils Bruin, Edgar Costa, Igor Dolgachev, Noam Elkies, Gavril Farkas, Sam Grushevsky, Avi Kulkarni, Elisa Lorenzo García, David Lubicz, Martin Raum, and Christophe Ritzenthaler for helpful conversations during the writing of this article. Furthermore, we thank Brendan Hassett, David Lehavi, Bjorn Poonen, and Jeroen Sijsling for reading a previous version of the manuscript and giving suggestions to improve the article.
@@ 2.1. The Prym canonical map
${S}$${C}$${\mathbb{P}^{2}}$${\Gamma_{\eta}\subseteq\mathbb{P}^{3}}$$\scriptstyle{b_{2}}$$\scriptstyle{b_{1}}$$\scriptstyle{\phi_{\eta}}$$\scriptstyle{c}$ (2.1)
@@ 2.2. Two linear maps
${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))\times_{\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}S^{2}\operatorname{H}^{0}(\Omega_{C})}$${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))}$${S^{2}\operatorname{H}^{0}(\Omega_{C})}$${\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}$ (2.2)
@@
$S^{2}\operatorname{H}^{0}(C,\Omega_{C}(\eta))\cong\operatorname{H}^{0}(\mathbb{P}^{2},\mathcal{O}(2))\overset{b_{1}^{*}}{\stackrel{{\scriptstyle\sim}}{{\rightarrow}}}\operatorname{H}^{0}(S,b_{1}^{*}\mathcal{O}(2))\cong\operatorname{H}^{0}\left(S,b_{2}^{*}\mathcal{O}(2)\otimes\mathcal{O}_{S}\left(-\sum_{i}N_{i}\right)\right)\,.$
@@
${C}$${S}$${\mathbb{P}^{2}}$${\mathbb{P}^{3}}$$\scriptstyle{\phi_{\eta}}$$\scriptstyle{b_{1}}$$\scriptstyle{b_{2}}$$\scriptstyle{c}$ (2.3)
@@
${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))}$${S^{2}\operatorname{H}^{0}(\Omega_{C})}$${\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}$$\scriptstyle{\psi}$
@@ 3.2. Finding the quadric
We will now construct an example over $\mathbb{F}_{37}$ and verify the required properties with the aid of Magma. (See the files Lemma-Section-3.2-proofs.m and construct-Lemma-Section-3.2-proofs.m in [HPS24b].) Consider the plane quartic $X\subset\mathbb{P}^{2}$ with equation
@@ 3.3. Finding a cubic
${S}$${\mathbb{P}(\operatorname{H}^{0}(C,\Omega_{C}(\eta)))\cong\mathbb{P}^{2}}$${\Gamma_{\eta}\hskip 5.0pt,}$$\scriptstyle{c}$
@@
$\begin{pmatrix}x\\ y\\ z\\ w\end{pmatrix}\mapsto\begin{pmatrix}33x^{2}+21xy+10xz+8xw+23y^{2}+30yz+13yw+34z^{2}+zw+11w^{2}\\ 6x^{2}+22xy+32xz+9xw+10y^{2}+7yz+31yw+16z^{2}+13zw\\ 8xy+15xz+22xw+28y^{2}+36yz+25yw+23z^{2}+23zw+2w^{2}\\ 35x^{2}+17xy+36xz+9xw+10yz+17yw+z^{2}+11zw+11w^{2}\\ 26x^{2}+17xy+26xz+17xw+8yz+29yw+2z^{2}+29zw+19w^{2}\\ 17xy+35xz+24xw+4y^{2}+34yz+18yw+4z^{2}+36zw+4w^{2}\end{pmatrix}\,.$
@@ 3.4. Main result
1
2Choose the two-torsion point given by $\eta=\frac{1}{2}\begin{bmatrix}0&0&0&0\\ 1&0&0&0\end{bmatrix}$.
3
4Compute the theta constants for the plane quartic $X$ via the Schottky-Jung relations
@@
Although our main result Theorem 3.11 requires the base field $k$ to be algebraically closed, with additional information one can recover equations over arithmetic base fields.
Over finite fields our method can be used to recover equations of genus $4$ curves from the knownledge of the algebraic theta constants. These are defined over a field extension of manageable size as the degree of this extension can be bounded using the fact that extensions of finite fields have cyclic Galois group.
By stark contrast, in the number field case the field of definition of the algebraic theta constants is typically a huge field extension of degree up to [Igu72][V, Theorem 4]
$[\operatorname{Sp}_{2g}(\mathbb{Z}):\Gamma(4,8)]=2^{g^{2}+g(2g+1)+2g}\prod_{i=1}^{g}(2^{2i}-1)\,.$
It is therefore imperative to use floating point approximations to the theta constants.
These can be computed analytically provided a small period matrix is known. In order to retrieve equations over a number field $K$, we need more information. If in addition, we are also given a big ($g\times 2g$) period matrix for $C$ with respect to a $K$-rational basis for $\operatorname{H}^{0}(C,\Omega_{C})$, we can recover equations for $C$ defined over $K$ as follows. First applying our theorem, we obtain approximate equations for $C$ with coefficients in $\operatorname{\mathbb{C}}$. Denote by $\widetilde{C}$ this complex curve, which is isomorphic to $C$ over $\operatorname{\mathbb{C}}$.
@@
When using the Schottky-Jung relations one takes a square root to compute the $\vartheta_{X}\big{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\big{]}$. This leads to a sign choice that has to be made correctly. For this one can use the discussion following [Gla80, Theorem 3.1] where Glass makes use of the quartic Riemann theta formulas to find a correct choice of signs. At this point, the assumption that the $\vartheta_{C}\big{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\big{]}$ are the theta constants of a genus $4$ curve is used. Indeed, the latter implies that the Schottky modular form vanishes. By construction of this modular form, this is equivalent to requiring that the $\vartheta_{X}\big{[}\begin{smallmatrix}a\\ b\end{smallmatrix}\big{]}$ satisfy the quartic Riemann theta formulas.
When the theta constants are given in an exact field, then taking square roots might lead to an auxilliary field extension. This field extension only appears in intermediate steps. The final result of Algorithm 3.1 consists of equations defined over the same field as the theta constants.
@@ 4. Examples
To remedy this, we call the command SomePrincipalPolarizations from the ModularCurves GitHub repository [S+23]. This produces several big period matrix candidates. Evaluating the Schottky modular form on the first of these yields a complex number with absolute value $6.2124\cdot 10^{-300}$ when calculated with precision 300, indicating that this principally polarized abelian variety is likely a Jacobian. Applying our method then produces the corresponding genus 4 curve, which, after a change of variable, is isomorphic to the curve $C$ in $\mathbb{P}^{3}$ given by the following equations.
@@ 5. Future work
To rule out CM fields containing primitive roots of unity, Moonen and Oort [MO11] defined the notion of a Weyl CM field. These are CM fields $K$ of degree $[K:\mathbb{Q}]=2g$ such that the Galois group of the splitting field is as large as possible, i.e., $(\mathbb{Z}/2)^{g}\rtimes S_{g}$. They put forth the conjecture that, for a given fixed $g\geqslant 4$, there are only finitely many curves of genus $g$ whose Jacobian has CM by a Weyl CM field.
@@ 5.2. Higher genus
In work in progress [HPS24a] we plan to generalize the Milne correspondence to arbitrary $g$. This would give a bijection between pairs of odd theta hyperplanes in $\mathbb{P}^{g-1}$ for $C$ and odd theta hyperplanes in $\mathbb{P}^{g-2}$ for the Prym variety $\operatorname{Prym}(C,\eta)$. Notice that in this context the Prym variety is no longer automatically a Jacobian. Nevertheless, one can define odd theta hyperplanes for arbitrary p.p.a.v.s; however, these will lack the geometric interpretation as multitangents of a canonically embedded curve (unless the p.p.a.v. happens to be a Jacobian).
@@
${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))\times_{\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}S^{2}\operatorname{H}^{0}(\Omega_{C})}$${S^{2}\operatorname{H}^{0}(\Omega_{C}(\eta))}$${S^{2}\operatorname{H}^{0}(\Omega_{C})}$${\operatorname{H}^{0}(\Omega_{C}^{\otimes 2})}$ (5.3)